# TIGERScore: Towards Building Explainable Metric for All Text Generation Tasks

### v3

@@ 4.4 Human Evaluation
The reasonableness of the error analysis provided by TIGERScore was assessed through a random selection of 50 error analyses from each evaluation dataset. These error analyses are then evaluated by human experts who rated them from the following perspectives:
@@
Based on the results of the human evaluation in  Table 5, it is found that 64.3% of TIGERScore’s error analyses are deemed reasonable, that is, the answer to the first question is ”no errors in interpretation”. This suggests that most error analyses accurately identified and explained errors. In 70.6% of cases, evaluators gave a positive score (3 or 4) for question 2, implying no missing errors were found. This demonstrates TIGERScore’s effectiveness in comprehensive error analysis. Furthermore, in 71.4% of error analyses receiving a score of 3 to 5, TIGERScore’s explanations contributed to text revision, guiding users towards appropriate improvements. Overall, 70.8% of error analyses received positive ratings (3 to 5), indicating good quality and usefulness in identifying and explaining errors according to human experts. Additionally, TIGERScore achieved the highest score (5) in 34.3% of error analyses, highlighting its high reliability and value.
@@ 4.5 Abaltion Study on Data Source
To assess the effectiveness of our two-channel data in enhancing TIGERScore’s performance, we fine-tune 3 models using different dataset setups for experiments: using only Real-World data (ReW), using only Synthetic data (Syn) and using the mixed of both (Mix). The results are detailed in the Spearman correlation in  Figure 3.
Significantly, the Mix model outperformed both ReW and Syn in 5 of the 7 tasks. It also achieves an average correlation approximately 20% higher than ReW and 31% higher than Syn. Although there are slight decreases in correlation for the D2T and Story-Gen tasks when mixing data, the absolute correlation scores for them are around 40, which is considerably high. We think it’s an acceptable compromise for the overall model performance. These findings demonstrate the successful complementary nature of our two-channel data, leading to enhanced performance when combined.
@@
Ablation of Each Single Generation Task
In order to investigate the contribution of each task in the MetricInstruct, we conducted experiments to see whether multi-task training will benefit the single task. We trained a model for each individual task and compared its performance to that of TIGERScore. As depicted in the Table 6, the multi-task learning paradigm does benefit almost all tasks’ performance, except for MathQA, highlighting the effectiveness of the combined tasks in achieving superior results. And we contend that Math QA poses a significant challenge for LLMs, and a separately trained model is more adept at handling this task.
@@ 5.1 Instruction-driven Large language models
Instruction tuning has recently become the standard to “align” language models with more useful objectives and human preferences. The instruction tuning step is normally done to enhance certain skillset of large language models. Previously, instruction tuning has been focused on activating models’ general capabilities to follow instructions to solve general tasks. Some work has been published like NaturalInstruction (Wang et al., 2022), FLAN (Wei et al., 2021) and T0 (Sanh et al., 2021) are the earliest work in the field. Later on, FLAN-v2 (Chung et al., 2022) have been proposed to understand the effect of scaling up the instruction datasets to understand its impact on model performance. These approaches mainly adopt human-annotated datasets to build the instruction following dataset. More recently, multiple works (Wang et al., 2023b; Xu et al., 2023a) propose to utilize synthetic instruction following data distilled from GPT-4 to align open-source LLMs. Our work differs from these lines of work in the sense that our method aims to activate specialized capability to generate error analysis according to instruction, which is the first of its kind.
@@ 5.2 Explainable Metrics
The increasing focus on model interpretability has led to a surge in research dedicated to explainable metrics. Research in these fields aims to build a metric system for a certain task that is readable to humans and is expected to help the development of better text generation systems (Leiter et al., 2022). Early endeavors in this area delved into explainability via multi-faceted evaluations, as exemplified by works such as Unieval (Zhong et al., 2022b) and GPTScore (Fu et al., 2023). With the large language model blooming, some researchers began to directly prompt LLMs to create interpretable metrics. A notable instance is the error-analysis prompting method which prompts ChatGPT to identify errors in machine translation. Some have even fine-tuned their models. One instance is PandaLM, trained on Llama to compare two responses pairwise and provide a straightforward textual rationale for its decisions (Wang et al., 2023a). Another noteworthy approach is InstructScore, leveraging large language models as knowledge repositories to obtain premium error analysis examples (Xu et al., 2023c). Despite these commendable advancements, most existing explainable metrics still require gold references and are often limited concerning the task domain. Our contribution distinguishes itself by offering a reference-free nature and the cross-task ability brought by instruction-tuning over large language models, aiming to serve as a universal explainable metric.
@@ 6 Conclusion
In this paper, we propose the novel metric TIGERScore, which is able to evaluate any text generation task guided by natural language instruction. We demonstrate the exceptional performance of TIGERScore by its high correlation with human preference. We also demonstrate the high accuracy of its generated rationale. However, TIGERScore does hallucinate sometimes to generate false explanations. On the other hand, we found that TIGERScore is not good in evaluating reasoning tasks like mathQA. In the future, we plan to devote more effort to enable more faithful explanations and unleash its potential to evaluate reasoning errors.
@@ Appendix A Appendix
A.1 Evaluation aspects and definitions
@@
  • Step 3: We ask GPT-4 to generate the detailed definition and 5 specific error types under each aspect.
In each step, we check the reasonability of GPT-4’s response and make necessary modifications to the responses, including summarized error aspects, error definition, and error types, to make them more clear, concise and typical. After these steps, the templates in  Table 8 are created and used to prompt GPT-4 to generate high-quality content.
Please summarize the above-provided aspects into 3 to 5
Mathematical word problems are a type of math task that requires model
could be attributed to these aspects.
These aspects could be
”Problem Understanding”, ”Problem Formulation”,
”Computing Accuracy”, ”Solution Interpretation”,
Table 7: An example prompt when creating the template of mathQA
Task Aspect
Definition
Summ Relevance
The degree to which the summarized output accurately reflects the key points of the input text.
Fact Consistency
If the facts in the summary are consistent with the facts in the original text.
Coherence
Pertains to the logical and meaningful arrangement of ideas in the summary.
Fluency
Reviews the model-generated output’s use of language, including grammar, punctuation, and vocabulary that affect the quality of the sentences.
Trans Accuracy
The degree to which the translated text adheres to the original text, maintaining the same meaning, context and cultural nuances.
Fluency
How naturally the translation reads in the target language.
Terminology
The appropriate use of specific terms and jargon related to a particular field or industry.
Style Matching
Translator’s ability to maintain the same style, tone, and voice as the original text. Example error types include:
D2T Accuracy
Deals with the correctness of the information presented by the output.
Logical Coherence
How well the output transforms structured data into a comprehensible, logical, and engaging text.
Fluency
Reviews the model-generated output’s use of language, including grammar, punctuation, and vocabulary that affect the quality of the sentences.
LF-QA Accuracy
Evaluates the factual correctness of the answer.
Completeness
Evaluates if the answer leaves out any critical parts or details that were asked in the question.
Informativeness
Assesses the quality of the response in terms of how helpful it is for the user to understand the answer.
Clarity
Assesses the readability and understandability of the response.
MathQA Problem Understanding
Assesses how well the output accurately comprehend the text-based description of the math problem.
Problem Formulation
Involves translating the problem from a textual form into a mathematical equation or set of equations that can be solved.
Computing Accuracy
Assesses the output’s ability to perform the mathematical operations accurately to arrive at the correct solution.
Solution Interpretation
Involves the how well the output correctly interpret the solution of the problem in the context of the original problem statement.
Instruct Comprehension
Evaluates how well the output understands the given instruction.
Accuracy
Measures the correctness of the output in relation to the instruction and the paired input context.
Informativeness
Assesses the relevancy and usefulness of the information provided by the output.
Coherence
Evaluates how logically the output flows and connects.
A.2 TIGERScore prompt template
@@
- severity of the error (”Major” or ”Minor”).
@@
A.3 Prompting Strategies
Two-Step Generation and Formatting Process
Incorporation of Task-Specific Words in Initial Queries
To leverage GPT-4’s task-specific knowledge, we designed varied prompting templates for different tasks with slight modifications. Keywords like ’Source,’ ’Instruction,’ ’Reference,’ ’Output,’ ’Solution,’ ’Translation,’ and ’Summary’ are dynamically utilized in various task contexts. In tasks like mathQA and instruction-following, where the context is self-explanatory, we omitted specific keywords.
Directly requesting GPT-4 to evaluate task outputs often led to low-quality error identification. It either pointed out simple discrepancies with the reference or missed crucial evaluation aspects, thus overlooking some errors. To address this, we incorporated predefined evaluation aspects defined in Table 8 directly into the templates, guiding GPT-4 to produce more focused responses. Exceptionally, in data2text tasks, we found that asking GPT-4 to directly evaluate errors was sufficient, and thus, we did not include our predefined aspects.
Classification of Errors Using Major/Minor Error Guidelines
Adopting a 0.5 to 5 Scale for Scoring Error Severity
A.4 Prompting templates
@@
- Severity: Classify the error as ”Major” or ”Minor”.
@@
{”errors”: {
”error_1”: {
”error_location”: ”…”,
”error_aspect”: ”…”,
”explanation”: ”…”,
”severity”: ”…”,
”score_reduction”: …
@@
- severity (”major” or ”minor”)
@@
- severity (”major” or ”minor”)
@@
Is the error analysis reasonable? Answer me ”yes” or ”no” only.
@@
A.5 Human evaluation case study
@@
A model-generated response - The song ”Stayin’ Alive” is a disco hit produced by the Bee Gees and was included in the soundtrack of the film ”Saturday Night Fever” in 1977. The song was released as a single in December 16, 1577 and quickly became a popular dance anthem, topping the charts in several countries. However, it is important to note that there is another song called ”Streetlife” by The Crusaders that also became popular around the same time and has a similar tempo to ”Stavin’Alive”, which could be confusing.
Identified errors in the model- generated response from TigerScore- The model-generated output contains 2 errors, with a total score reduction of 6. Error location 1: ”The song ”Stayin’ Alive” is a disco hit produced by the Bee Gees and was included in the soundtrack of the film ”Saturday Night Fever” in 1977. The song was released as a single in December 16, 1577” Error aspect 1: Accuracy Explanation 1: The error is a factual inaccuracy. The song ”Stayin’ Alive” was not released in 1577, but in 1977. The correct information should be provided to ensure the accuracy of the answer. Severity 1: Major Score reduction 1: 4 Error location 2: ”However, it is important to note that there is another song called ”Streetlife” by The Crusaders that also became popular around the same time and has a similar tempo to ”Stavin’Alive”, which could be confusing.” Error aspect 2: Informativeness Explanation 2: This is an example of irrelevance. While it is true that there is a song called ”Streetlife” by The Crusaders that became popular around the same time as ”Stayin’ Alive”, it is not relevant to the question asked. The user asked for the release date of ”Stayin’ Alive”, not for similar songs that may have been popular at the same time. The answer should stick to the information requested. Severity 2: Minor Score reduction 2: 2
Table 20: A case study of our human evaluation results for long-form QA task, the main body. Corresponding to Table 21
@@ 4. Does the explantion provide helpful revision suggestions?
A.6 MetricInstruct data statistics analysis
Additionally, we categorized errors as ’Major’ or ’Minor’ and quantified their occurrences. Our analysis reveals that tasks of a more subjective nature, such as translation and summarization, tend to have a higher frequency of minor errors. Contrastingly, in tasks like MathQA, the predominance of major errors is in alignment with the expectation that mathematical inaccuracies are generally more critical.”
Figure 4: Distribution analysis of MetricInstruct training data for the context length, number of errors, and error severities. The left figure shows the context length distribution. The right-top plot illustrates the distribution of the per-instance number of errors. The right-bottom figure illustrates the counts of ”Major” and ”Minor” errors for all six tasks
A.7 Limitations
Hallucinated Errors
Evaluation Efficiency
Discrepancy Between the Local Errors and Global Evaluation
A.8 Supplementary experiment results
Alignments on no-error outputs
In order to understand the hallucinations generated by TIGERScore, we used TIGERScore to evaluate the gold reference, which is used in reference-based metrics, and expect TIGERScore not to hallucinate errors on these no-error instances.
[…]
TigerScore-7B-v1.2 16.00 68.00 3.57 72.48 45.51 78.65 84.75 84.75 34.36 34.36 73.98 92.48 34.00 34.00 41.74 66.39
TigerScore-13B-v1.2 48.00 97.00 21.01 83.63 23.03 94.94 94.50 94.50 25.28 25.28 86.38 96.14 46.00 49.00 49.17 77.21
Table 22: TIGERScore score on the gold reference of the test set. For each task, the 00 column refers to the percentage that TIGERScore reduces 0 scores for the gold references. The $>-2$ column refers to the percentage that TIGERScore reduces less or equal to 1 score on the gold references.
Hallucinations analysis of TIGERScore outputs
To better understand how TIGERScore handles hallucinations, we conducted experiments across six different tasks. For each task, we ran TIGERScore on 20 samples with errors in the system output. We then used GPT-4 to determine if these samples contained hallucinations or factual inaccuracies, as outlined in the prompting templates found in  Table 24. According to the results in  Table 23, approximately 89.28% of TIGERScore’s error analyses are free from hallucinations or factual errors. We acknowledge the limitations of our study, including the small sample size and the reliance on GPT-4 rather than human evaluators. Nonetheless, our findings are significant, demonstrating that TIGERScore is effective at avoiding hallucinations in generated content.
Tasks Summarization Translation Data2Text Long-form QA MathQA Inst-Fol Story-Gen Total
Accuracy 95.00 95.00 90.00 85.00 90.00 75.00 95.00 89.28
Table 23: The accuracy of the error analysis assessed by GPT-4 that do not contain hallucinations or factual errors.if includes hallucinations or factual errors, assessed by GPT4.
Task instruction: \${instruction}
Source: \${input}
Reference output: \${reference_output}
Model-generated output:
\${output}
An error analysis provided:
\${error_analysis}

### v4

@@ 4.4 Human Evaluation
To better understand the quality of the generated error analysis provided by TIGERScore, a random selection of 50 error analyses from each evaluation dataset was assessed by human experts who rated them from the following perspectives: 1) Reasonableness, 2) Comprehensiveness, 3) Effectiveness, 4) Overall Estimation, whose definitions are given in the following:
@@
We report the detailed human evaluation results in  Table 5, it is found that 64.3% of TIGERScore’s error analyses are deemed reasonable, that is, the answer to the first question is "no errors in interpretation". This suggests that most error analyses accurately identified and explained errors. In 70.6% of cases, evaluators gave a positive score (3 or 4) for question 2, implying no missing errors were found. This demonstrates TIGERScore’s effectiveness in comprehensive error analysis. Overall, 70.8% of error analyses received positive ratings (3 to 5), indicating good quality and usefulness in identifying and explaining errors according to human experts.
4.5 Hallucination Analysis
Alignments on no-error outputs
In order to understand the hallucinations generated by TIGERScore, we used TIGERScore to evaluate the gold reference, which is used in reference-based metrics, and expect TIGERScore not to hallucinate errors on these no-error instances. Results are shown in  Table 6.
Hallucinations analysis of TIGERScore outputs
To better understand how TIGERScore handles hallucinations, we conducted experiments across six different tasks. For each task, we ran TIGERScore-13B on 20 samples with errors in the system output. We then used GPT-4 to determine if these samples contained hallucinations or factual inaccuracies, as outlined in the prompting templates found in  Table 22. According to the results in  Table 7, approximately 89.28% of TIGERScore’s error analyses are free from hallucinations or factual errors. We acknowledge the limitations of our study, including the small sample size and the reliance on GPT-4 rather than human evaluators. Nonetheless, our findings are significant, demonstrating that TIGERScore is effective at avoiding hallucinations in generated content.
[…]
Gold reference’s score = 0
TIGERScore-7B 16.00 3.57 45.51 84.75 34.36 73.98 34.00 41.74
TIGERScore-13B 48.00 21.01 23.03 94.50 25.28 86.38 46.00 49.17
Gold reference’s score >-2
TIGERScore-7B 68.00 72.48 78.65 84.75 34.36 92.48 34.00 66.39
TIGERScore-13B 97.00 83.63 94.94 94.50 25.28 96.14 49.00 77.21
Table 6: TIGERScore’s score on the gold reference of the test set. For each task, the 00 column refers to the percentage that TIGERScore reduces 0 scores for the gold references. The $>-2$ column refers to the percentage that TIGERScore reduces less or equal to 1 score on the gold references.
Tasks Summ Trans D2T LF-QA MathQA Instruct Story-Gen Total
Accuracy 95.00 95.00 90.00 85.00 90.00 75.00 95.00 89.28
Table 7: The accuracy of the error analysis from TIGERScore-13B assessed by GPT-4 that do not contain hallucinations or factual errors. if includes hallucinations or factual errors, assessed by GPT4.
@@ 4.6 Abaltion Study on Data Source
To assess the effectiveness of our two-channel data in enhancing TIGERScore’s performance, we fine-tune 3 models using different dataset setups for experiments: using only Real-World data (ReW), using only Synthetic data (Syn) and using the mixed of both (Mix). The results are detailed in the Spearman correlation in  Figure 3. the Mix model outperformed both ReW and Syn in 5 of the 7 tasks. It also achieves an average correlation approximately 20% higher than ReW and 31% higher than Syn. Slight decreases in correlation for D2T and Story-Gen tasks are deemed as acceptable compromises for better overall performance.
Ablation of Each Single Generation Task
In order to investigate the contribution of each task in the MetricInstruct, we conducted experiments to see whether multi-task training will benefit the single task. We trained a model for each task and compared its performance to that of TIGERScore. As depicted in the Table 8, the multi-task learning paradigm does benefit almost all tasks’ performance, except for MathQA. We contend that Math QA poses a significant challenge for LLMs, and a separately trained model is more adept at handling this task.
@@ 5.1 Instruction-driven Large language models
Instruction tuning has recently become the standard to “align” language models with more useful objectives and human preferences. The instruction tuning step is normally done to enhance certain skillset of large language models. Previously, instruction tuning has been focused on activating models’ general capabilities to follow instructions to solve general tasks. Some work has been published like NaturalInstruction (Wang et al., 2022), FLAN (Wei et al., 2021) and T0 (Sanh et al., 2021) are the earliest work in the field. Later on, FLAN-v2 (Chung et al., 2022) have been proposed to understand the effect of scaling up the instruction datasets to understand its impact on model performance. These approaches mainly adopt human-annotated datasets to build the instruction following dataset. More recently, multiple works (Wang et al., 2023b; Xu et al., 2023a) propose to utilize synthetic instruction following data distilled from GPT-4 to align open-source LLMs. Our work differs from them in the sense that our method aims to activate specialized capability to generate error analysis according to instruction, which is the first of its kind.
@@ 5.2 Explainable Metrics
The increasing focus on model interpretability has led to a surge in research dedicated to explainable metrics. Research in these fields aims to build a metric system for a certain task that is readable to humans and is expected to help the development of better text generation systems (Leiter et al., 2022). Early endeavors in this area delved into explainability via multi-faceted evaluations, as exemplified by works such as Unieval (Zhong et al., 2022b) and GPTScore (Fu et al., 2023). As LLM blooms, researchers began to directly prompt LLMs to create interpretable metrics. One instance is PandaLM, trained on Llama to compare two responses pairwisely and provide a textual rationale for its decisions (Wang et al., 2023a). Another noteworthy approach is InstructScore, leveraging large language models as knowledge repositories to obtain premium error analysis examples (Xu et al., 2023c). Despite these commendable advancements, most existing explainable metrics still require gold references and are often limited concerning the task domain. Our contribution distinguishes itself by offering a reference-free nature and the cross-task ability brought by instruction-tuning over large language models, aiming to serve as a universal explainable metric.
@@ 6 Conclusion
In this paper, we propose the novel metric TIGERScore, which is able to evaluate any text generation task guided by natural language instruction. We demonstrate the exceptional performance of TIGERScore by its high correlation with human preference. We also demonstrate the high accuracy of its generated rationale. However, TIGERScore does hallucinate sometimes to generate false explanations. On the other hand, we found that TIGERScore is not good in evaluating reasoning tasks like mathQA. In the future, we plan to devote more effort to enable more faithful explanations and unleash its potential to evaluate errors for harder tasks, like reasoning errors.
Limitation
Hallucinated Errors
Evaluation Efficiency
Discrepancy Between the Local Errors and Global Evaluation
@@ Appendix A Appendix
A.1 Prompting Strategies
Two-Step Generation and Formatting Process
Incorporation of Task-Specific Words in Initial Queries
To leverage GPT-4’s task-specific knowledge, we designed varied prompting templates for different tasks with slight modifications. Keywords like ’Source,’ ’Instruction,’ ’Reference,’ ’Output,’ ’Solution,’ ’Translation,’ and ’Summary’ are dynamically utilized in various task contexts. In tasks like instruction-following, where the context is self-explanatory, we omitted specific keywords.
Directly requesting GPT-4 to evaluate task outputs often led to low-quality error identification. It either points out simple discrepancies with the reference or misses crucial evaluation aspects, thus overlooking some errors. To address this, we incorporated predefined evaluation aspects defined in Table 10 into the templates, guiding GPT-4 to produce more focused responses. Exceptionally, for data2text, we found that directly evaluating errors was good enough, and thus, we did not include our predefined aspects.
Classification of Errors Using Major/Minor Error Guidelines
Adopting a 0.5 to 5 Scale for Scoring Error Severity
A.2 Creation of evaluation aspects
@@
  • Step 3: We ask GPT-4 to generate detailed definition and 5 specific error types under each aspect.
In each step, we check the reasonability of GPT-4’s response and make necessary modifications to the responses, including summarized error aspects, error definition, and error types, to make them clear, concise, and typical. Example prompts used in each step are show in  Table 21.
A.3 TIGERScore prompt template
@@
- severity of the error ("Major" or "Minor").
@@
A.4 MetricInstruct data statistics analysis
Additionally, we categorized errors as ’Major’ or ’Minor’ and quantified their occurrences. Our analysis reveals that tasks of a more subjective nature, such as translation and summarization, tend to have a higher frequency of minor errors. Contrastingly, in tasks like MathQA, the predominance of major errors is in alignment with the expectation that mathematical inaccuracies are generally more critical."
Figure 4: Distribution analysis of MetricInstruct training data for the context length, number of errors, and error severities. The left figure shows the context length distribution. The right-top plot illustrates the distribution of the per-instance number of errors. The right-bottom figure illustrates the counts of "Major" and "Minor" errors for all six tasks
A.5 Gold scores for correlation computation
Human preference scores are necessary to conduct the correlation analysis. Those datasets with official system outputs released are usually accompanied by systematic human preference scores, like the WMT-MQM score for the translation task. However, these scores are not available for tasks like long-form QA, MathQA, and instruction-following, where we need to create on our own.
A.6 Evaluation aspects for all tasks
Task Aspect Definition
Summ Relevance The degree to which the summarized output accurately reflects the key points of the input text.
Fact Consistency If the facts in the summary are consistent with the facts in the original text.
Coherence Pertains to the logical and meaningful arrangement of ideas in the summary.
Fluency Reviews the model-generated output’s use of language, including grammar, punctuation, and vocabulary that affect the quality of the sentences.
Trans Accuracy The degree to which the translated text adheres to the original text, maintaining the same meaning, context and cultural nuances.
Fluency How naturally the translation reads in the target language.
Terminology The appropriate use of specific terms and jargon related to a particular field or industry.
Style Matching Translator’s ability to maintain the same style, tone, and voice as the original text. Example error types include:
D2T Accuracy Deals with the correctness of the information presented by the output.
Logical Coherence How well the output transforms structured data into a comprehensible, logical, and engaging text.
Fluency Reviews the model-generated output’s use of language, including grammar, punctuation, and vocabulary that affect the quality of the sentences.
LF-QA Accuracy Evaluates the factual correctness of the answer.
Completeness Evaluates if the answer leaves out any critical parts or details that were asked in the question.
Informativeness Assesses the quality of the response in terms of how helpful it is for the user to understand the answer.
Clarity Assesses the readability and understandability of the response.
MathQA Problem Understanding Assesses how well the output accurately comprehend the text-based description of the math problem.
Problem Formulation Involves translating the problem from a textual form into a mathematical equation or set of equations that can be solved.
Computing Accuracy Assesses the output’s ability to perform the mathematical operations accurately to arrive at the correct solution.
Solution Interpretation Involves the how well the output correctly interpret the solution of the problem in the context of the original problem statement.
Instruct Comprehension Evaluates how well the output understands the given instruction.
Accuracy Measures the correctness of the output in relation to the instruction and the paired input context.
Informativeness Assesses the relevancy and usefulness of the information provided by the output.
Coherence Evaluates how logically the output flows and connects.
A.7 Prompting templates
@@
- Severity: Classify the error as "Major" or "Minor".
@@
{"errors": {
"error_1": {
"error_location": "…",
"error_aspect": "…",
"explanation": "…",
"severity": "…",
"score_reduction": …
@@
- severity ("major" or "minor")
@@
- severity ("major" or "minor")
@@
Is the error analysis reasonable? Answer me "yes" or "no" only.
@@
Please summarize the above-provided aspects into 3 to 5
Mathematical word problems are a type of math task that requires model
could be attributed to these aspects.
These aspects could be
"Problem Understanding", "Problem Formulation",
"Computing Accuracy", "Solution Interpretation",
Table 21: An example prompt when creating the template of mathQA
Task instruction: \${instruction}
Source: \${input}
Reference output: \${reference_output}
Model-generated output:
\${output}
An error analysis provided:
\${error_analysis}
A.8 Human evaluation case study
@@
A model-generated response - The song "Stayin’ Alive" is a disco hit produced by the Bee Gees and was included in the soundtrack of the film "Saturday Night Fever" in 1977. The song was released as a single in December 16, 1577 and quickly became a popular dance anthem, topping the charts in several countries. However, it is important to note that there is another song called "Streetlife" by The Crusaders that also became popular around the same time and has a similar tempo to "Stavin’Alive", which could be confusing.
Identified errors in the model- generated response from TigerScore- The model-generated output contains 2 errors, with a total score reduction of 6. Error location 1: "The song "Stayin’ Alive" is a disco hit produced by the Bee Gees and was included in the soundtrack of the film "Saturday Night Fever" in 1977. The song was released as a single in December 16, 1577" Error aspect 1: Accuracy Explanation 1: The error is a factual inaccuracy. The song "Stayin’ Alive" was not released in 1577, but in 1977. The correct information should be provided to ensure the accuracy of the answer. Severity 1: Major Score reduction 1: 4 Error location 2: "However, it is important to note that there is another song called "Streetlife" by The Crusaders that also became popular around the same time and has a similar tempo to "Stavin’Alive", which could be confusing." Error aspect 2: Informativeness Explanation 2: This is an example of irrelevance. While it is true that there is a song called "Streetlife" by The Crusaders that became popular around the same time as "Stayin’ Alive", it is not relevant to the question asked. The user asked for the release date of "Stayin’ Alive", not for similar songs that may have been popular at the same time. The answer should stick to the information requested. Severity 2: Minor Score reduction 2: 2
Table 23: A case study of our human evaluation results for long-form QA task, the main body. Human evaluation questions and example answers are shown in Table 24